The sum of John's age and Jorge's age is 56. Ten years ago John was twice as old as Jorge. How old is Jorge now?
Answer by rfer(16322)

let john =2x+10=34
let jorge =x+10=22
(2x+10)+(x+10)=56
3x+20=56
3x=36
x=12